Prev Next

SW&IT Meeting Minutes 2019-07-22

Attendees: Andrew, David, Igor, Matthieu, Petr, Stephen, Xianyu, Yang

Outlook

Labor Reporting

Report auto-generated every, Do not edit this table, use entries in smartsheets.

News from management

Previous Action Items

Open topics

  • techtalk - C++11 (PK)

Team reports

MB

  • vacation
next
  • vacation

DE

  • Work on other sensor platform for Peter
  • Get into Vault backup script
    • Powershell was written by a team of crack sadists
    • Seriously, why did they not just rip off Unix shells?
    • Adding considerable functionality and robustness
      • Check to see if there are process locks, kill process if there are, alert if we can't
      • Wrap offsite backup into same script
Next:
  • Track down updates on colo project approval
  • More time in the Powershell dungeon
  • UITS wants a switch upgrade in our server room 6 Aug 0400-0700

YZ

  • Major rework of the LBC skyflat wiki program (hopefully to release a new version this week).
  • Restarted to work on AO-Cravat web interface polishing.
  • Working with Olga on several LBC/MODS related issues:
    • LBC/MODS alarm scripts and several problems.
    • MODS computer disk usage.
    • Science image archival.
  • Need to work on merging sciops-lbcfpia and sciops-lbcwrs on github into one repo.

PK

  • Ice3.7 - pull request, adding commands to it2tcs (to mimics IRC)
  • TCS presentations
  • Turned off PMC, MCS & OVMS telemetry
next
  • Ice3.7, finish it2tcs
  • AGW1 troubleshooting
  • AOS & GCS

AC

  • IT 6405/7705 - Removed obsolete DDS parameters from wda but symptoms still persist. Perhaps increase sleep time or make two separate while true loops.
  • FACSUM - repeated successful build on cisrv01 and deployed on the test NUC
  • updated MagExposure script on azcamserver
next:
  • IT 7604 - AGW telemetry
    • gather requirements, define streams of data to log, hdf5, etc.

SH

  • archive firewall
  • laptop for Tyler
  • test NUC (fixes to kickstart)
  • LBTI DNS
  • Carl setup
  • AGW SOUL network Alessandro
  • tcs1, tcs2, tcs-test upgraded, tcs3 removed from loop.
  • ECS telemetry restart after upgrade
  • read OVMS-TT document
  • talked to Chris Johnson about colocating non-critical machines
next:
  • swap out hall NUC
  • ks another 507 machine, install in six-head location
  • github John and Steve's stuff
  • update switch firmware as I can

XZ

Last week
  • read and sent comments to OVMS-TT
  • Workedd on DX ASM on the summit
    • removed DX ASM from its hub
    • hooked and powered and a baseline was acquired
    • magnet test routine was tested successfully
    • completed the testing of the detached magnet routine and run the routine over all the actuators
    • the thin shell (TS3) was removed from DX ASM
Next week:
  • work on detached magnet data
  • read the new OVMS-TT version

IS

Last week
  • OT
    • Finished OT TPE submenu work
    • Went through Joe's git fixes related to curved slits and added them to the current release.
    • Upgraded Dithering Pattern form to avoid confusion between pattern and the application of a scale factor and offsets with no scale factor applied.

This week:
  • OT
    • Make sure slopes in slanted slits on MOS mask have the proper sign
    • remove camera defaulting based on grating for LUCI
    • Make a test build
    • Start working on OT related MODS issues

CH

  • Monnit telemetry
    • POC: Peter Gray
    • Goal: liberate the sensor tm from imonnit.com via REST API.
    • Modified PHP script to dump all sensor data to console.
    • Peter opted for me to complete integration HDF5 rather than using JSON files.
    • Translating to Python for simpler integration with HDF5 API.
  • Alarm Handling System (AHS, aka CHAOS?)
    • POC: Matthieu Bec, Petr Kubanek
    • Goal: overhaul/consolidate disparate alerting tools.
    • Installed existing ALH demo on workstation
    • Reviewing copies of known alert scripts.
    • Will poll other PIs to seek out unknown alerting scripts.

Topic revision: r16 - 05 Aug 2019, MatthieuBec
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback